OpenStudy (karatechopper):

is -1/2x a linear equation? is -1/2x a linear equation? @Mathematics

myininaya (myininaya):

\[y=-\frac{1}{2}x \text{ is linear}\]

OpenStudy (anonymous):

yes it is if you mean y=-x/2 the its a linear line

myininaya (myininaya):

\[y=-\frac{1}{2x} \text{ is \not linear }\]

OpenStudy (karatechopper):

k thts what i thought cuz some one else is confusing me on my other question

myininaya (myininaya):

anything in this form or anything that can be written in this form y=mx+b is a linear equation

OpenStudy (karatechopper):

i kno

OpenStudy (karatechopper):

can u graph tht for me? i need to check if i am right on my graph

myininaya (myininaya):

y=-x/2?

myininaya (myininaya):

y=mx+b m=slope=rise/run=-1/2 b=y-intercept=0 so the line goes through the origin so count down 1 and go over right twice for your next pt connect the dots and then you are done

OpenStudy (karatechopper):

i did tht but draw what the line looks like so i am sure

OpenStudy (karatechopper):

is it like this?:

OpenStudy (karatechopper):

myin?

OpenStudy (anonymous):

if its y= -x/2 then it is indeed more or less like that yes

OpenStudy (karatechopper):

yay!!! thnx!:)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-x%2F2

OpenStudy (karatechopper):

cool!